FIND THE QUOTIENTS;
1 364.19/10
2 479.23/100

Answer by checkley71(8403)

364.19/10=36.419 ANSWER.
479.23/100=4.7923 ANSWER.
----------------------------------------
WHEN DIVIDING BY 10,100,1000,10000 ETC. JUST MOVE THE DECIMAL TO THE LEFT THE NUMBER OF ZEROES IN THE DENOMINATOR.
---------------------------------------------------------------------
WHEN MULTIPLYING A NUMBER BY 10,100,1000,10000 THEN MOVE THE DECIMAL THE NUMBER OF ZEROES TO THE RIGHT.
45.263*100=4526.3
.37415*1000=374.15
87.4293*10=874.293